[Brahms-l] Shift report 20040401 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Apr 01 2004 - 00:05:58 EST
------------------------------------------------------------
Date: Thu Apr  1 00:05:57 2004, Shift: 16:00-24:00
Supervisor: ykl, with: 
Log book#15 pages 103 - 103

Shift Summary:
------------------------------------------------------------
<pre>
1. 4.25 hours of good beam. Then RF problem. Beams
are still out at this time.

2. MRS at 90 deg, 6KG B.
   FS at 8 deg, 1/(sqrt 2)^3 A

    
            Tr.3  Tr.4
Run 11513   14K   26K;
Run 11520   18K   32K;
Run 11521   15K   27K
  
Run 11523    4K    7K;
(run11523 is a special run with low scale down
to increase statistics for minimum bias)

Run 11524-11525 pedestal runs.

3. DAQ problem, connection problem with fb3 and
fb4 partitions;
Pii4 and pii12 problems. solved by rebooting.
</pre>
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 1 00:06:12 2004

This archive was generated by hypermail 2.1.8 : Thu Apr 01 2004 - 00:06:35 EST